A coloured solution is placed within a spectrometer. At 554 nm the solution shows an absorbance of 0.69. Calculate the percentage of light that is being absorbed.


1
Expert's answer
2021-05-08T23:33:51-0400


Wavelength, "\lambda=554nm,λ=554nm, Absorbance = 0.69=0.69"


The relation between Absorbance and transmittance is-


"A=-log(\dfrac{T}{100})

\\[9pt]

0.69=-log(\dfrac{T}{100})
\\[9pt]


\dfrac{T}{1000}=10^{-0.69}"


"T=0.2041\times 100=20.41" %


Percentage of light absorbed


"= 100-T=100-20.41=79.59" %